find three consecutive odd numbers that the sum of the last two bigger numbers is 85 more thean the first
Answer by ankor@dixie-net.com(22740)

find three consecutive odd numbers
x, (x+2), (x+4)
:
that the sum of the last two bigger numbers is 85 more than the first
(x+2) + (x+4) = x + 85
2x + 6 = x + 85
2x - x = 85 - 6
x = 79
:
79, 81, 83 are the numbers
:
Check:
81 + 83 = 79 + 85